OpenStudy (anonymous):

This question wants me to put the information into a little graph thing (which ill draw below) and its a bit confusing for me XD help?? Eric has a gift card to his favorite restaurant. Every Tuesday he has the same meal at the restaurant. What is the equation that represents Eric’s gift card balance? Record any fractional answers in decimal form. (y-100)=m(x-b)

OpenStudy (anonymous):



OpenStudy (anonymous):

if you notice....when the week doubles the card amount looses $25.00 so double thr week again (make it week 10) and subtract 25 from the 50

OpenStudy (anonymous):

yah she is right

OpenStudy (anonymous):

Oh no im not trying to find the next row im trying to solve this equation y - 100 = m(x - b)

OpenStudy (anonymous):

@katti

OpenStudy (anonymous):

I think the solved equation is y - 100 = 12.5( x - 4 )

OpenStudy (anonymous):

no its y - 100 = -12.5( x - 4 )

OpenStudy (anonymous):

your right